OpenStudy (anonymous):

what is c^-1 (x) given c (x) =x^3-5 ?

OpenStudy (anonymous):

it is the same as the previous one try to solve it and show me ur work =)

OpenStudy (amistre64):

the inverse of c(x); c(x)=y replace "x" with "y" and solve for "x"

OpenStudy (amistre64):

x = y^3 - 5 x+5 = y^3 3root(x+5) = y c^-1(x) = 3root(x+5)